OpenStudy (anonymous):

Given that c (ax + by - cz) = b (ax - by + cz) = a ( -ax + by + cz) = 2 then ax + by + cz/ab + bc + ca = ?

OpenStudy (perl):

Let's unpack the first statement. c (ax + by - cz) = b (ax - by + cz) = a ( -ax + by + cz) = 2

OpenStudy (perl):

If we look at the first two statements c (ax + by - cz) = 2 b (ax - by + cz) = 2 It follows that (ax + by - cz) = 2/c (ax - by + cz) = 2/b now you can add these two equations to eliminate variables x and y 2ax = 2/c + 2/b

OpenStudy (perl):

can you solve that for x ?

OpenStudy (anonymous):

Yes. How do we exactly approach such questions? Is there any trick?

OpenStudy (perl):

What I wrote is clear so far?

OpenStudy (anonymous):

x = 1/ac + 1/ab

OpenStudy (anonymous):

Okay

OpenStudy (perl):

The approach I used is common to solving linear equations, try to find a way to eliminate a variable. Well not exactly this approach, after some work

OpenStudy (perl):

lets be more careful with solving for x

OpenStudy (perl):

2ax = 2/c + 2/b you divided out 2 correctly ax = 1/c + 1/b first simplify 1/c + 1/b ax = ( b + c ) / (bc) now divide by a x = ( b + c ) / (abc)

OpenStudy (perl):

$$ \large \frac{1}{b} + \frac{1}{c} = \frac{1}{b}\cdot \frac {c}{c} + \frac{1}{c} \cdot \frac{b}{b}= \frac{c}{bc} + \frac{b}{bc} = \frac{b+c}{bc}$$

OpenStudy (anonymous):

Thank you @perl